# Second-order Optimization of Gaussian Splats with Importance Sampling

### v1

@@ 6 Discussion and Conclusion
Limitations. While our method is computationally efficient for a lower number of Gaussians, its advantages diminish as the number increases, which may limit scalability in highly complex scenes. Exploring importance sampling within the local neighborhood of the Gaussians could be an interesting research direction to address this limitation. Moreover, we employ mean squared error instead of mean absolute error and only used a diagonal approximation of the original SSIM loss, unlike the original Adam-based optimization for 3D Gaussian Splatting. This discrepancy in loss formulations may affect convergence behavior, but future work could explore re-weighting strategies of residuals to obtain different norms.
By leveraging the inherent sparsity of the Jacobian matrix and integrating a GPU-parallelized conjugate gradient solver, our method significantly reduces both memory consumption and computational overhead. Our novel view and pixel-wise importance sampling further enhance efficiency, enabling rapid convergence by decreasing the per-step overhead. Additionally, our heuristic for learning rate selection eliminates the need for costly line search procedures, further accelerating training. Our approach achieves up to a 6× speedup over Adam, particularly excelling in scenarios with a low number of Gaussians. Overall, our results highlight the potential of second-order methods in accelerating optimization for 3D Gaussian Splatting. We anticipate that future work will refine these techniques, particularly in handling larger Gaussian counts and incorporating additional perceptual loss terms, further advancing the efficiency and quality of Gaussian-based scene representations.
Supplementary Material
@@ Appendix A Details of CUDA Implementation
We follow the parallelization pattern applied in 3DGS [21], dividing images into $16\times 16$ blocks, and launching threads for each pixel. Our main difference from 3DGS is that our block size is significantly less than $16\times 16=256$, due to pixel sampling. We set the number of pixels sampled per block as multiples of $32$, which allows us to do fast warp level reductions when needed. Now, we describe the Jacobian-vector product kernels in detail:
Jv Kernel. This kernel is similar to backward pass of 3DGS, and can be split into two kernel calls. We launch a thread for sampled columns (which corresponds to pixels) of $\mathbf{J}^{\top}$, and loop over the Gaussians. However, when the sample size is small, i.e 32, only 32 threads (or one warp per block) are launched. The limited number of threads does not fully utilize the GPU’s resources, leading to inefficiencies. To mitigate, we launch additional threads along the Gaussian direction. Each thread processes part of the Gaussian loop, skipping most calculations except for four dependent variables (transmittance and accumulated RGB values). These values are simply recalculated by each thread. Warp level reductions then finalize the dot product.
Jv Kernel. This kernel also uses per pixel parallelization and computes the dot product. However, it cannot be split into two smaller parts as the $\mathbf{J}^{\top}\mathbf{v}$ kernel. Therefore, we utilize the shared memory more by bringing additional Gaussian parameters. Warp-level reductions are not needed; each thread independently computes and writes its local sum.
DiagJJ Kernel. Similarly, this kernel cannot be split, so we again leverage shared memory to hold more data. As we compute derivatives, we square and accumulate them using warp-level reductions.
For the rest of the conjugate gradient algorithm, we follow the implementation described in [52]. This version requires fewer kernel calls when compared to a naive implementation and has been successfully applied in other works [16, 20, 57]. For the details, please see the respective papers.
Table 5: Comparisons on Tanks &\&& Temples scenes. All optimizers use MSE loss.
@@
3DGS-Adam Train 0.321 0.670 20.12 130.0
LM-IS (Ours) 0.302 0.694 20.16 145.28
3DGS-Adam Truck 0.185 0.780 23.44 128.18
LM-IS (Ours) 0.203 0.772 22.93 119.83
3DGS-Adam Avg. 0.253 0.725 21.78 129.10
LM-IS (Ours) 0.253 0.733 21.55 132.56
Appendix B Results Per Scene
We share the per-scene scores obtained in NeRF synthetic dataset [35] in Table 6. In this experiment, all of the optimizers use MSE loss.
@@ Appendix B Results Per Scene
Method Scene LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\downarrow$ Time (s)
3DGS-RMSprop Chair 0.162 0.865 24.96 54.11
3DGS-Adam 0.187 0.834 22.09 60.22
Taming3DGS-Adam 0.145 0.876 25.75 61.32
LM-IS (Ours) 0.147 0.874 25.39 13.53
3DGS-RMSprop Drums 0.185 0.834 22.05 52.73
3DGS-Adam 0.164 0.864 24.92 55.36
Taming3DGS-Adam 0.166 0.845 22.57 52.82
LM-IS (Ours) 0.166 0.845 22.30 12.26
3DGS-RMSprop Ficus 0.090 0.893 25.79 56.21
3DGS-Adam 0.093 0.890 25.60 56.92
Taming3DGS-Adam 0.078 0.903 26.51 60.74
LM-IS (Ours) 0.094 0.891 25.55 10.01
3DGS-RMSprop Hotdog 0.111 0.923 29.65 53.72
3DGS-Adam 0.111 0.924 29.78 57.85
Taming3DGS-Adam 0.099 0.929 30.34 60.99
LM-IS (Ours) 0.088 0.933 30.55 13.29
3DGS-RMSprop Lego 0.196 0.814 24.34 53.55
3DGS-Adam 0.201 0.809 24.23 57.73
Taming3DGS-Adam 0.184 0.821 24.82 59.97
LM-IS (Ours) 0.180 0.827 24.88 13.88
3DGS-RMSprop Materials 0.182 0.826 23.95 51.48
3DGS-Adam 0.183 0.827 23.95 56.11
Taming3DGS-Adam 0.167 0.829 24.22 62.11
LM-IS (Ours) 0.158 0.862 24.21 11.73
3DGS-RMSprop Mic 0.104 0.922 28.14 50.48
3DGS-Adam 0.108 0.920 28.09 56.02
Taming3DGS-Adam 0.100 0.925 28.64 63.34
LM-IS (Ours) 0.093 0.930 28.10 11.80
3DGS-RMSprop Ship 0.238 0.770 25.28 63.03
3DGS-Adam 0.274 0.769 25.27 67.41
Taming3DGS-Adam 0.238 0.767 25.44 64.49
LM-IS (Ours) 0.225 0.784 25.59 15.57
3DGS-RMSprop Average 0.159 0.856 25.52 54.42
3DGS-Adam 0.162 0.55 25.49 58.46
Taming3DGS-Adam 0.147 0.862 26.01 60.73
LM-IS (Ours) 0.144 0.867 25.82 12.76
We also provide experiments with SSIM loss, which is approximated diagonally, as mentioned in the main paper. The original SSIM loss propagates gradients through the local neighbors, which is not efficient with per-pixel parallelization scheme of our $\mathbf{Jv}$ kernel. In Table 7, we show that our optimizer still outperforms the competing methods even with the diagonal approximation. All optimizers use a weighted average of MSE and SSIM losses, with weights $1.0$ and $0.2$, respectively.
@@
Table 7: Comparisons of different methods on the synthetic NeRF dataset. All optimizers use a weighted average of MSE and SSIM loss, with weights $1.0$ and $0.2$ respectively.
@@
Method Scene LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\downarrow$ Time (s)
3DGS-RMSprop Chair 0.155 0.878 24.88 85.70
3DGS-Adam 0.157 0.878 24.88 87.42
Taming3DGS-Adam 0.138 0.890 25.66 86.58
LM-IS (LM-IS (Ours)) 0.135 0.888 25.88 38.93
3DGS-RMSprop Drums 0.190 0.853 21.52 83.96
3DGS-Adam 0.190 0.853 21.60 86.46
Taming3DGS-Adam 0.159 0.872 22.38 86.82
LM-IS (LM-IS (Ours)) 0.152 0.870 22.69 36.68
3DGS-RMSprop Ficus 0.089 0.901 25.34 84.59
3DGS-Adam 0.090 0.901 25.41 83.50
Taming3DGS-Adam 0.074 0.913 26.28 87.88
LM-IS (LM-IS (Ours)) 0.082 0.907 25.85 33.08
3DGS-RMSprop Hotdog 0.098 0.937 29.47 86.06
3DGS-Adam 0.096 0.938 29.64 85.20
Taming3DGS-Adam 0.087 0.942 30.29 84.79
LM-IS (LM-IS (Ours)) 0.078 0.945 31.28 36.92
3DGS-RMSprop Lego 0.189 0.835 24.15 83.55
3DGS-Adam 0.192 0.836 24.170 86.41
Taming3DGS-Adam 0.175 0.847 24.68 85.47
LM-IS (Ours) 0.165 0.846 25.27 43.77
3DGS-RMSprop Materials 0.175 0.864 23.84 81.72
3DGS-Adam 0.176 0.863 23.80 85.35
Taming3DGS-Adam 0.159 0.870 24.12 85.55
LM-IS (Ours) 0.146 0.871 24.28 41.16
3DGS-RMSprop Mic 0.103 0.936 27.90 81.66
3DGS-Adam 0.109 0.933 27.75 80.23
Taming3DGS-Adam 0.090 0.941 28.30 84.91
LM-IS (Ours) 0.083 0.942 28.48 34.26
3DGS-RMSprop Ship 0.227 0.804 22.77 86.02
3DGS-Adam 0.225 0.806 24.90 88.33
Taming3DGS-Adam 0.214 0.811 25.15 85.38
LM-IS (Ours) 0.219 0.806 25.71 46.40
3DGS-RMSprop Average 0.153 0.876 25.23 84.16
3DGS-Adam 0.154 0.876 25.27 85.37
Taming3DGS-Adam 0.137 0.886 25.86 85.93
LM-IS (Ours) 0.132 0.885 26.18 38.90
Appendix C Results in Real-world Datasets
We share our experiment results on the Tanks &\&& Temples dataset [24]. In this dataset, the initial number of Gaussians is larger than $100{,}000$. Although the performance of our optimizer drops, we can still compete with the Adam optimizer as shown in Table 5 and Fig 6. In this dataset, Adam optimizer is run for $10{,}000$ iterations, while ours is run for $130$ iterations. We use MSE loss for both of the optimizers.
Appendix D Example of Learning Rate Schedule
Figure 5: Our learning rate heuristic can assign both high and low learning rates dynamically, resulting in stable and rapid convergence.
Our learning rate scheduler can assign both low and high learning rates, while providing stable training. We share an example of learning rates obtained in Lego scene in Fig. 5. Note that we cap the maximum learning rate to $0.5$ in our experiments.
Truck Scene Train Scene
3DGS-Adam
Refer to caption Refer to caption Refer to caption Refer to caption
LM-IS (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: We show qualitative results on the real-world dataset. Our second-order optimizer reaches a similar quality in a similar amount of time.

### v2

@@ 6 Discussion and Conclusion
Limitations. We use mean squared error instead of mean absolute error and only a diagonal approximation of the original SSIM loss, unlike the original Adam-based optimization for 3D Gaussian Splatting. This discrepancy in loss formulations may affect convergence behavior, but future work could explore re-weighting strategies of residuals to obtain different norms.
By leveraging the inherent sparsity of the Jacobian matrix and integrating a GPU-parallelized conjugate gradient solver, our method significantly reduces the computational overhead. Our novel view and pixel-wise sampling further enhance efficiency, enabling rapid convergence by decreasing the per-step overhead. Additionally, our dynamic learning rate scheduler eliminates the need for costly line search procedures, further accelerating training. Our approach achieves up to $5\times$ speedup over Adam, particularly excelling in scenarios with low number of Gaussians. Overall, our results highlight the potential of second-order methods in accelerating optimization for 3D Gaussian splatting. We anticipate that future work will refine these techniques, particularly incorporating additional loss terms, further advancing the efficiency and quality of Gaussian-based scene representations.
@@ Appendix A Details of CUDA Implementation
Algorithm 2 Forward Pass - 3DGS
Input: Sorted Gaussians projected on each pixel: $\mathcal{G}$.
Output: Rendered RGB color C.
1:for in parallel pixel $p$ in ALL pixels do
2:  float T = 1.0 $\triangleright$ Initial Transmittance
3:  Vec3 C = 0.0 $\triangleright$ Accumulated Color
4:  for $g$ in $\mathcal{G}$[$p$] do
5:    // Read Gaussian Properties
6:    Vec2 mean2D = $g$.getMean2D()
7:    Mat2x2 invCov2D = $g$.getInverseCov2D()
8:    float opac = $g$.getOpacity()
9:    Vec3 color = $g$.getColor()
10:
11:    // Compute the exponential fall-off
12:    Vec2 d = mean2D - pp.getPixelCenter()
13:    float power = $\frac{-1}{2}\cdot d\cdot\text{invCov2D}\cdot d^{T}$
14:    float alpha = $\text{opac}\times e^{\text{power}}$
15:
16:    // Update output color and transmittance
@@
18:    T = T $\times\,\text{(1-alpha)}$
19:  end for
20:  return C
21:end for
Algorithm 3 Automatic Forward Differentiation - LM-RS
Input: Sorted Gaussians projected on each pixel: $\mathcal{G}$.
Output: Rendered RGB color C.
1:for in parallel pixel $p$ in SAMPLED pixels do
2:  DualFloat T = (1.0, 0.0) $\triangleright$ Initial Transmittance
3:  DualVec3 C = (0.0, 0.0) $\triangleright$ Accumulated Color
4:  for $g$ in $\mathcal{G}$[$p$] do
5:    // Read Gaussian Properties
6:    DualVec2 mean2D = $g$.getMean2D()
7:    DualMat2x2 invCov2D = $g$.getInverseCov2D()
8:    DualFloat opac = $g$.getOpacity()
9:    DualVec3 color = $g$.getColor()
10:
11:    // Compute the exponential fall-off
12:    DualVec2 d = mean2D - pp.getPixelCenter()
13:    DualFloat power = $\frac{-1}{2}\cdot d\cdot\text{invCov2D}\cdot d^{T}$
14:    DualFloat alpha = $\text{opac}\times e^{\text{power}}$
15:
16:    // Update output color and transmittance
[…]
18:    T = T $\times\,\text{(1-alpha)}$
19:  end for
20:  return C.prime
21:end for
22:
23:class DualFloat
24:  float real
25:  float prime
[…]
27:    return DualFloat(real $\times$ other.real,
28:    real $\times$ other.prime + prime $\times$ other.real) $\triangleright$ Product Rule
29:  end function
30:  // Similarly, overload other operators (+, - , /, $e^{x}$, …)
31:end class
32:// Similarly, define DualVec and DualMatrix classes.
We follow the parallelization pattern applied in 3DGS [3dgs], dividing images into $16\times 16$ blocks, and launching a thread for each pixel. Our main difference from 3DGS is that our block size is significantly less than $16\times 16=256$, due to pixel sampling. In our implementation, we only support sample sizes that are multiples of $32$, which allows us to do fast warp-level reductions when needed. Now, we describe our CUDA kernels in detail:
Jv Kernel. This kernel is similar to the backward pass of 3DGS and is executed in two kernel calls using backward mode automatic differentiation. The first kernel computes and forwards the derivatives associated with splatting operations to the Gaussians, while the second kernel computes the derivatives related to Gaussian preprocessing. The main difference with 3DGS is that the number of threads per block is significantly reduced due to sampling, enabling us to perform faster computations. We use warp-level reductions to collect the partial dot products when needed.
Jv Kernel. This kernel implements forward‐mode automatic differentiation via the dual‐number data structure. We overload each operator so that, in a single pass, it performs both the real‐value computation and its derivative. This approach is more efficient than using backward-mode automatic differentiation when the Jacobian-vector product is needed [baydin2018automatic].
Conceptually, this kernel works similarly to the forward pass of 3DGS, but enhanced to carry the derivative information. Therefore, unlike Jv kernel, it first carries out Gaussian preprocessing, then implements splatting. Since we should additionally carry the derivative information, the shared memory usage is doubled compared to the forward pass of 3DGS. For a comparison between the forward pass of 3DGS and our forward-mode automatic differentiation, see Alg. 2 and Alg. 3. As it can be observed from the algorithm, there is no need for warp reductions in this kernel.
Diag ($J$) Kernel. To our current knowledge, there is no easy way to split this kernel into two calls. Therefore, we create a single kernel that squares and accumulates derivatives using warp-level reductions.
Since the remaining part of the preconditioned conjugate gradient algorithm requires only vector-related computations, it can be efficiently implemented in PyTorch.
Appendix B More Results in Synthetic Dataset
We compare our optimizer against Adam [adam], RMSprop [rmsprop], and SGD with momentum [sgd_momentum], using their respective implementations in PyTorch. The momentum term of SGD and second-moment factor of RMSprop is set to $0.99$. For SGD, we used the learning rates $0.16$ for the location, $0.2$ for color and $0.1$ for other parameters. We also apply a learning rate schedule for the mean value parameters, as suggested in 3DGS. The gradients are computed using either the vanilla 3DGS or Taming 3DGS [taming]. All baselines are run for 10,00010{,}000 iterations. We provide the per-scene scores obtained in the NeRF synthetic dataset [nerf] in Table 6. Please also see the supplementary video, comparing Adam and our optimizer in real-time. Additionally, we provide visual results of competing methods in Fig. 6.
Results with SSIM Loss: We also conduct experiments using SSIM loss, which is approximated diagonally, as mentioned in the limitations section of the main paper. The usual SSIM loss propagates gradients through the local neighbors, which is hard to implement with the parallelization scheme of our $\mathbf{Diag(J^{\top}J)}$ kernel. In Table 7, we show that our optimizer still outperforms the competing methods even with the diagonal approximation. However, to compensate for this discrepancy, we ran our method for $300$ iterations instead of $200$ iterations. The weights are set to $1.0$ and $0.2$, for MSE and SSIM terms respectively. All other hyperparameters remain identical to those specified in the main paper.
Figure 6: We visualize the test set performance of baseline methods evaluated on the Lego scene. Our method achieves comparable metrics in significantly less time. All optimizers use the MSE loss function.
Appendix C More Results in Real-world Scenes
We report per‐scene results for the Mip‐NeRF360 Indoor, DeepBlending, and Tanks & Temples datasets in Tab. 8, Tab. 9, and Tab. 10, respectively. All optimizers are trained with the mean‐squared error (MSE) loss. Notably, our method outperforms both the Adam optimizer and 3DGS‐LM on the Mip‐NeRF360 Indoor and DeepBlending datasets, and beats 3DGS‐LM on Tanks & Temples.
Memory Requirements: As it can be observed from the tables, our method uses more memory than Adam but significantly less memory than 3DGS-LM. When updating $n$ parameters, Adam optimizer scales with the number of parameters and uses $O(3n)$ space to store the gradient and two momentum estimates. Our matrix-free solver also scales with the number of parameters; however, we need 8 vectors to run the PCG algorithm, making the space complexity $O(8n)$. This means both Adam and our optimizer have the same $O(n)$ memory complexity, and our memory requirement differs by only a constant factor. On the other hand, 3DGS-LM scales with both the number of pixels ($p$) and the number of parameters, as they need to cache the Jacobian and save the PCG-related vectors, resulting in $O(np)$ space complexity.
Results with SSIM Loss: Both 3DGS-LM and our method employ the diagonal approximation of the SSIM loss. In Tab. 11, we compare both methods and show that our optimizer converges faster than 3DGS-LM (w/o Adam). However, in this setting, the diagonal approximation becomes overly restrictive, and 3DGS-LM outperforms our method by leveraging the full SSIM loss thanks to the Adam optimizer.
Appendix D Details of Dynamic Learning Rate Scheduler
Our learning rate scheduler can assign both low and high learning rates, while providing stable training. We share the obtained learning rates in Mip-NeRF360 scenes in Fig. 7. Note that, in all experiments, we cap the maximum learning rate to $0.2$ to avoid overshooting, and we use a fixed learning rate $0.05$ in the first $10$ iterations.
Figure 7: Our dynamic learning rate scheduler can assign both high and low learning rates, resulting in stable and rapid convergence.
Appendix E Alternative Probability Distributions for Residual Sampling
We experimented with different pixel (residual) sampling strategies to improve the performance. The following candidates are considered for comparison:
$\mathbf{q}_{\textbf{uniform}}$: This distribution assigns equal probability $\frac{1}{M}$ to every pixel $p$, where M is the total number of pixels. This is the distribution used in the main paper.
$\mathbf{q}_{\text{uniform}}(p)=\frac{1}{M}$ (16)
$\mathbf{q}_{\textbf{residual}}$: This distribution weights pixel $p$ according to its residual error. After rendering the Gaussians to obtain the image $\hat{I}$, we compute the per‐pixel residuals against the ground truth image $I$. We then normalize these residuals using the softmax function:
$r_{p}$ $=\hat{I}_{p}-I_{p}$ (17)
$\mathbf{q}_{\text{residual}}(p)$ $=Softmax(\left|r_{p}\right|)=\frac{e^{\left|r_{p}\right|}}{\sum_{k}^{M}e^{\left|r_{k}\right|}}$ (18)
$\mathbf{q}_{\textbf{Gaussian}}$: This distribution weights pixel $p$ by the number of Gaussians contributing to it, which we denote as $G_{p}$. During the forward pass, we count the number of Gaussians that contribute to pixels’ rendering, omitting those that lie beyond the visibility range (i.e., low transmittance) or are transparent. This weight is then normalized by the total number of Gaussians used across all $M$ pixels.
$\mathbf{q}_{\text{Gaussian}}(p)=\frac{G_{p}}{\sum_{k}^{M}G_{k}}$ (19)
We compare the performance of the these distributions in Tab. 5. While other importance sampling methods yield improvements in certain scenes, their average behavior closely matched that of $\mathbf{q}_{\text{uniform}}$. Accordingly, we adopt $\mathbf{q}_{\text{uniform}}$ in our paper and conduct all experiments with it. The similarity in the results could be because of the fixed sample size per tile (32) in every experiment. Our implementation restricts sample counts to multiples of 32, matching the CUDA warp size to enable efficient warp-level reductions. Investigating lower sample sizes while avoiding warp divergence can be an interesting direction to explore.
Table 5: This table compares the alternative distribution functions for residual sampling. The results are obtained by averaging across Mip-NeRF360 indoor scenes.
Distribution LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
$\mathbf{q}_{\text{Gaussian}}$ 0.160 0.859 28.65
$\mathbf{q}_{\text{residual}}$ 0.161 0.857 28.49
$\mathbf{q}_{\textbf{uniform}}$ (Ours) 0.161 0.859 28.59
Appendix F Densification of Gaussians
3DGS proposes a densification algorithm to increase the number of Gaussians, which is disabled in this work. The reason is that densification algorithms typically rely on statistics collected over many iterations and complement the dynamics of the Adam optimizer. However, second‐order optimizers require far fewer iterations, making it difficult to integrate such algorithms directly. Because of similar reasons, densification is also disabled in 3DGS-LM. We believe that combining the second-order optimizer with one of the densification strategies would be an interesting research direction.
@@ Appendix F Densification of Gaussians
Method Scene LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
3DGS-RMSprop Chair 0.162 0.865 24.96 54
3DGS-Adam 0.187 0.834 22.09 60
Taming3DGS-Adam 0.145 0.876 25.75 61
LM-RS (Ours) 0.138 0.882 25.96 11
3DGS-RMSprop Drums 0.185 0.834 22.05 52
3DGS-Adam 0.164 0.864 24.92 55
Taming3DGS-Adam 0.166 0.845 22.57 52
LM-RS (Ours) 0.159 0.855 22.48 10
3DGS-RMSprop Ficus 0.090 0.893 25.79 56
3DGS-Adam 0.093 0.890 25.60 56
Taming3DGS-Adam 0.078 0.903 26.51 60
LM-RS (Ours) 0.086 0.897 25.76 10
3DGS-RMSprop Hotdog 0.111 0.923 29.65 53
3DGS-Adam 0.111 0.924 29.78 57
Taming3DGS-Adam 0.099 0.929 30.34 60
LM-RS (Ours) 0.085 0.936 31.04 13
3DGS-RMSprop Lego 0.196 0.814 24.34 53
3DGS-Adam 0.201 0.809 24.23 57
Taming3DGS-Adam 0.184 0.821 24.82 59
LM-RS (Ours) 0.173 0.833 25.14 12
3DGS-RMSprop Materials 0.182 0.826 23.95 51
3DGS-Adam 0.183 0.827 23.95 56
Taming3DGS-Adam 0.167 0.829 24.22 62
LM-RS (Ours) 0.152 0.859 24.54 10
3DGS-RMSprop Mic 0.104 0.922 28.14 50
3DGS-Adam 0.108 0.920 28.09 56
Taming3DGS-Adam 0.100 0.925 28.64 63
LM-RS (Ours) 0.090 0.930 28.56 25
3DGS-RMSprop Ship 0.238 0.770 25.28 63
3DGS-Adam 0.274 0.769 25.27 67
Taming3DGS-Adam 0.238 0.767 25.44 64
LM-RS (Ours) 0.228 0.782 25.62 13
3DGS-RMSprop Average 0.159 0.856 25.52 54
3DGS-Adam 0.162 0.855 25.49 58
Taming3DGS-Adam 0.147 0.862 26.01 60
LM-RS (Ours) 0.139 0.872 26.14 13
@@
Table 7: Comparisons of different methods on the NeRF synthetic dataset. All optimizers use a weighted average of MSE and SSIM loss, with coefficients of $1.0$ and $0.2$ respectively. Our method provides $\approx 4\times$ speedup over 3DGS-Adam.
@@
Method Scene LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
3DGS-RMSprop Chair 0.155 0.878 24.88 85
3DGS-Adam 0.157 0.878 24.88 87
Taming3DGS-Adam 0.138 0.890 25.66 86
LM-RS (Ours) 0.131 0.892 26.18 22
3DGS-RMSprop Drums 0.190 0.853 21.52 83
3DGS-Adam 0.190 0.853 21.60 86
Taming3DGS-Adam 0.159 0.872 22.38 86
LM-RS (Ours) 0.153 0.870 22.66 21
3DGS-RMSprop Ficus 0.089 0.901 25.34 84
3DGS-Adam 0.090 0.901 25.41 83
Taming3DGS-Adam 0.074 0.913 26.28 87
LM-RS (Ours) 0.077 0.911 26.27 20
3DGS-RMSprop Hotdog 0.098 0.937 29.47 86
3DGS-Adam 0.096 0.938 29.64 85
Taming3DGS-Adam 0.087 0.942 30.29 84
LM-RS (Ours) 0.076 0.946 31.46 25
3DGS-RMSprop Lego 0.189 0.835 24.15 83
3DGS-Adam 0.192 0.836 24.170 86
Taming3DGS-Adam 0.175 0.847 24.68 85
LM-RS (Ours) 0.163 0.847 25.31 23
3DGS-RMSprop Materials 0.175 0.864 23.84 81
3DGS-Adam 0.176 0.863 23.80 85
Taming3DGS-Adam 0.159 0.870 24.12 85
LM-RS (Ours) 0.140 0.874 24.70 20
3DGS-RMSprop Mic 0.103 0.936 27.90 81
3DGS-Adam 0.109 0.933 27.75 80
Taming3DGS-Adam 0.090 0.941 28.30 84
LM-RS (Ours) 0.086 0.942 28.65 26
3DGS-RMSprop Ship 0.227 0.804 22.77 86
3DGS-Adam 0.225 0.806 24.90 88
Taming3DGS-Adam 0.214 0.811 25.15 85
LM-RS (Ours) 0.226 0.800 25.53 25
3DGS-RMSprop Average 0.153 0.876 25.23 84
3DGS-Adam 0.154 0.876 25.27 85
Taming3DGS-Adam 0.137 0.886 25.86 85
LM-RS (Ours) 0.132 0.885 26.34 23
Table 8: Comparisons of competing methods in Mip-NeRF360 Indoor scenes. Our method provides $\approx 1.3\times$ speedup over Adam, $\approx 2.1\times$ over 3DGS-LM, and $\approx 6.2\times$ over 3DGS-LM (w/o Adam). For 3DGS-LM number of iterations is calculated as: Number of First Order Iterations + Number of Second Order Iterations
[…]
3DGS counter 155767 0.190 0.822 26.63 253 5.02 10,000
3DGS-LM (w/o Adam) - - 25.42 1332 - 30
3DGS-LM - - 26.71 495 - 10,000 + 10
LM-RS (Ours) 0.189 0.829 26.83 183 15.03 200
3DGS kitchen 241367 0.136 0.857 28.03 277 5.83 10,000
3DGS-LM (w/o Adam) - - 26.42 1294 - 30
3DGS-LM - - 28.15 536 - 10,000 + 10
LM-RS (Ours) 0.134 0.864 28.61 180 15.12 200
3DGS room 112627 0.176 0.858 29.52 240 6.37 10,000
3DGS-LM (w/o Adam) - - 27.85 1087 - 30
3DGS-LM - - 29.51 271 - 10,000 + 5
LM-RS (Ours) 0.177 0.861 29.61 190 14.21 280
3DGS bonsai 206613 0.141 0.877 28.65 240 6.04 10,000
3DGS-LM (w/o Adam) - - 27.17 1031 - 30
3DGS-LM - - 28.85 290 - 10,000 + 5
LM-RS (Ours) 0.142 0.881 29.29 208 14.66 210
3DGS Average 0.161 0.854 28.21 252.5 5.81
3DGS-LM (w/o Adam) - - 26.72 1186 $\approx$ 50
3DGS-LM - - 28.31 398 $\approx$ 50
LM-RS (Ours) 0.161 0.859 28.59 190.25 14.75
Table 9: Comparisons of competing methods in DeepBlending scenes. Our method provides $\approx 1.5\times$ speedup over Adam.
[…]
3DGS playroom 37005 0.205 0.850 26.90 200 3.05 10,000
LM-RS (Ours) 0.197 0.854 27.61 122 6.42 400
3DGS drjohnson 80861 0.275 0.840 26.74 218 4.11 10,000
LM-RS (Ours) 0.274 0.841 26.83 152 8.35 400
3DGS Average 0.240 0.846 26.82 209 3.58
LM-RS (Ours) 0.236 0.848 27.22 137 7.39
Table 10: Comparisons of competing methods in Tanks & Temples scenes. Our method provides $\approx 1.5\times$ speedup over 3DGS-LM and $\approx 3.4\times$ over 3DGS-LM (w/o Adam) optimizer. For 3DGS-LM, the number of iterations is calculated as: Number of First Order Iterations + Number of Second Order Iterations
[…]
3DGS train 182686 0.321 0.671 20.22 120 2.12 10,000
3DGS-LM (w/o Adam) - - 20.09 454 - 15
3DGS-LM - - 20.74 172 - 10,000 + 5
LM-RS (Ours) 0.298 0.687 20.22 117 6.20 130
3DGS truck 136029 0.186 0.779 23.45 117 1.78 10,000
3DGS-LM (w/o Adam) - - 23.05 291 - 10
3DGS-LM - - 23.87 171 - 10,000 + 5
LM-RS (Ours) 0.208 0.774 23.43 100 5.43 130
3DGS Average 0.254 0.725 21.84 118.5 1.95
3DGS-LM (w/o Adam) - - 21.57 372.5 $\approx$ 40
3DGS-LM - - 22.31 171 $\approx$ 40
LM-RS (Ours) 0.253 0.731 21.83 108.5 5.82
Table 11: Comparisons of competing methods in Tanks & Temples scenes. Our method provides $\approx 6.5\times$ speedup over 3DGS-LM (w/o Adam) while using $6.8\times$ less memory. The results are obtained using a weight of 1.0 for the MSE loss and 0.25 for the SSIM loss. For 3DGS-LM, the number of iterations is calculated as: Number of First Order Iterations + Number of Second Order Iterations.
[…]
3DGS-LM (w/o Adam) train 182686 19.62 604 - 30
3DGS-LM 20.52 174 - 10,000 + 5
LM-RS (Ours) 19.99 118 6.20 130
3DGS-LM (w/o Adam) truck 136029 23.05 828 - 20
3DGS-LM 23.61 173 - 10,000 + 5
LM-RS (Ours) 23.13 103 5.43 130
3DGS-LM (w/o Adam) 21.34 716 $\approx$ 40
3DGS-LM 22.07 173 $\approx$ 40
LM-RS (Ours) 21.56 110 5.82